# On Card guessing after a single shelf shuffle

### v1

@@ 1.1. Shuffling techniques and card guessing games
The shuffling of cards and associated card guessing games have a long and rich history in mathematics at the intersection of combinatorics, probability theory and statistics, dating back to work of Markov [24] and Poincare [27]. Famous shuffling techniques include the riffle shuffle [1, 9, 13], also called dove-tail shuffling, as well as the shelf shuffle [4, 5]; we refer the reader to [28] for references concerning other shuffling techniques such as the Thorpe shuffle and the clumpy shuffle and to [18] for the Monge shuffle. Additionally, we refer to the book of Diaconis and Fulman [7] for many more references. The different card shuffling techniques and their analysis are not only of purely mathematical interest. There are applications to the analysis of clinical trials [2, 10], fraud detection related to extra-sensory perceptions [6], guessing so-called Zener Cards [25], as well as relations to tea tasting and the design of statistical experiments [11, 26]. In this work we analyze the shelf shuffle, as introduced by Diaconis, Fulman and Holmes [8]. Its importance stems from card shuffling machines called shelf-shuffler, which play an important role in the theory of card guessing, as sch machines have direct practical applications to casino gambling and manufacturing. A standard shuffling machine employs ten shelves. In a shuffling machine the cards are drawn from the top of the deck. Each card is placed, uniformly at random, on a shelf. The decision to place the card on the top or bottom of the existing pile on the shelf is also done randomly. The shelves are then removed and the cards are pressed together, forming a shuffled deck. Clay [4] analyzed properties of a single shelf shuffle, including a card guessing game with complete feedback. An ordered deck of $n$ cards labeled $1$ up to $n$ is shelf-shuffled exactly one time. Then, one after the other a single card is drawn from the shuffled deck. The guesser wants to correctly guess the label of the drawn cards to maximze the number of correct guesses. The person gets full feedback after the guess: the card is always shown until no cards remain. Let $X_{n}$ denote the number of correct guesses in the full feedback card guessing game after a single shelf shuffle. Clay obtained the expected value of $X_{n}$ and showed that ${\mathbb{E}}(X_{n})=\frac{3}{4}\cdot n$. We note here in passing that similar questions recently received very much attention and mention a few corresponding works for the riffle shuffle [3, 19, 20, 21, 23] and also refer to the recent work of Tripathi [29] and also to Clay [5] for related problems. The purpose of this work is to extend the analysis of $X_{n}$ and to obtain the distribution, the variance, as well as a central limit theorem. Additionally, we introduce a biased, asymmetric shelf shuffle with parameter $p\in(0,1)$ and show how to extend the analysis to this case. Finally, motivated by other card guessing games [22] we also discuss a refined analysis of $X_{n}$: correct guesses are either pure luck $L_{n}$, as they occur with probability one-half, or they are certified correct $C_{n}$ with probability one. Thus, we obtain a refinement of $X_{n}$ into
$X_{n}=L_{n}+C_{n},$
and we study the joint distribution of both random variables.
@@ 1.2. Notation
2. Shelf Shuffle
We begin by recalling the definition of the shelf shuffle.
@@ 2. Shelf Shuffle
Given an ordered deck of $n$ cards labelled one up to $n$. Cards are drawn from the bottom of the deck and placed in a pile. When a card is drawn, it is placed at the top or at the bottom of the pile with probability 1/2.
@@ 3.2. Distributional analysis
$S(z,v)=\frac{zv+\frac{\frac{z^{2}v}{2}}{1-\frac{zv}{2}}}{1-\frac{zv}{2}-\frac{\frac{z}{2}}{1-\frac{zv}{2}}+\frac{z}{2}}=\frac{2zv\big(2-(1-v)z\big)}{4-4vz+(v^{2}-v)z}.$
@@ 3.3. Expectation, variance and limit law
$\rho_{1}(v)=\frac{2(1-\sqrt{v})}{v(v-1)},\quad\rho_{2}(v)=\frac{2(1+\sqrt{v})}{v(v-1)},$
@@ 4.1. Asymmetric shelf shuffle
For $p\geq 1/2$ we can carry out analysis as before, as the optimal strategy stays the same. The distributional equation changes only slightly, as stated below, and the generating functions approach works directly exactly the same:
with initial values $X_{1}=1$ and $X_{0}=0$ and
For $p<\frac{1}{2}$ the structure is asymptotically similar, but the optimal strategy changes, as we do not guess the number one, as the guess depends on the value of $p$. We have to guess the label(s) corresponding to the maximum of $\{p(1-p)^{i-1},\,1\leq i\leq n-1\}$ and $(1-p)^{n-1}$. However, the optimal strategy may not be unique.
@@
$p=\frac{1}{6}\left(2^{2/3}\sqrt[3]{\sqrt{93}-9}-2^{2/3}\sqrt[3]{9+\sqrt{93}}+6\right)\approx 0.31767<1/2$
@@
For $p\in[\frac{1}{2},1]$ we proceed as in the case $p=\frac{1}{2}$: we guess the card labelled one as the first guess and then always the card immediately following the previously shown card in all other positions. Once one of card n1n-1 or $n$ is first shown, the player should guess the remaining cards in descending order, and is guaranteed to get these correct. For $p\in[0,\frac{1}{2})$ we guess the label $\operatorname{maxP}_{n}$, corresponding to the maximum of of $\{p(1-p)^{i-1},\,1\leq i\leq n-1\}$ and $(1-p)^{n-1}$. Then, after the card, say labelled $j$ is shown to the guesser, we proceed in a similar way, but with a reduced set of indices $\{1,\dots,n-j\}$ and guess the label corresponding to the maximum of $\{p(1-p)^{i-1},\,1\leq i\leq n-j-1\}$ and $(1-p)^{n-j}$, shifted of course by $j$.
@@
with initial values $X_{1}=1$ and $X_{0}=0$ and the random variable $\operatorname{FC}$ as stated in (1). Here, $\mathbb{I}_{n}(\operatorname{FC}_{n}=n^{\ast})$ denotes the indicator of the event that the guess is correct and $X^{\ast}$ denotes an independent copy of $X$ and the random variables appearing in the decomposition are mutually independent.
@@ 4.2. Refined analysis of the ordinary shelf shuffle
One can apply higher dimensional quasi-power theorem analogs of Heuberger and Kropf [14], leading to a central limit theorem, with variance-covariance governed by(11). We omit the details.
@@ 5. Outlook and summary
A corresponding analysis of the guessing game without feedback seems to be more involved, insofar as even the optimal strategy is unclear. A spectral analysis of the position matrix for the asymmetric model is doable. Of interest is also the transition for $p\to 0$ or $p\to 1$ in the asymmetric model and to look for transitions from normal to say Poisson or degenerated.

### v2

@@ 1.1. Shuffling techniques and card guessing games
The shuffling of cards and associated card guessing games have a long and rich history in mathematics at the intersection of combinatorics, probability theory and statistics, dating back to work of Markov [24] and Poincare [27]. Famous shuffling techniques include the riffle shuffle [1, 9, 13], also called dove-tail shuffling, as well as the shelf shuffle [4, 5]; we refer the reader to [28] for references concerning other shuffling techniques such as the Thorpe shuffle and the clumpy shuffle and to [18] for the Monge shuffle. Additionally, we refer to the book of Diaconis and Fulman [7] for many more references. The different card shuffling techniques and their analysis are not only of purely mathematical interest. There are applications to the analysis of clinical trials [2, 10], fraud detection related to extra-sensory perceptions [6], guessing so-called Zener Cards [25], as well as relations to tea tasting and the design of statistical experiments [11, 26]. In this work we analyze the shelf shuffle, as introduced by Diaconis, Fulman and Holmes [8]. Its importance stems from card shuffling machines called shelf-shuffler, which play an important role in the theory of card guessing, as sch machines have direct practical applications to casino gambling and manufacturing. A standard shuffling machine employs ten shelves. In a shuffling machine the cards are drawn from the top of the deck. Each card is placed, uniformly at random, on a shelf. The decision to place the card on the top or bottom of the existing pile on the shelf is also done randomly. The shelves are then removed and the cards are pressed together, forming a shuffled deck. Clay [4] analyzed properties of a single shelf shuffle, including a card guessing game with complete feedback. An ordered deck of $n$ cards labeled $1$ up to $n$ is shelf-shuffled exactly one time. Then, one after the other a single card is drawn from the shuffled deck. The guesser wants to correctly guess the label of the drawn cards to maximze the number of correct guesses. The person gets full feedback after the guess: the card is always shown until no cards remain. Let $X_{n}$ denote the number of correct guesses in the full feedback card guessing game after a single shelf shuffle. Clay obtained the expected value of $X_{n}$ and showed that ${\mathbb{E}}(X_{n})=\frac{3}{4}\cdot n$. We note here in passing that similar questions recently received very much attention and mention a few corresponding works for the riffle shuffle [3, 19, 20, 21, 23] and also refer to the recent work of Tripathi [29] and of Clay [5] for related problems. The purpose of this work is to extend the analysis of $X_{n}$ and to obtain the distribution, the variance, as well as a central limit theorem for $X_{n}$. Additionally, we introduce a biased, asymmetric shelf shuffle with parameter $p\in(0,1)$ and show how to extend the analysis to this case and also uncovering a phase transition for $p\to 1$. Finally, motivated by other card guessing games [22] we also discuss a refined analysis of $X_{n}$: correct guesses are either pure luck $L_{n}$, as they occur with probability one-half, or they are certified correct guesses $C_{n}$ with probability one. Thus, the random variable $X_{n}$ decomposes into the random variables $L_{n}$ and $C_{n}$, counting pure luck and certified correct guesses, respectively:
$X_{n}=L_{n}+C_{n}.$
We study the joint distribution of both random variables, shedding more light into the distribution of $X_{n}$; in particular, we prove that exactly $n/2$ guesses are expected to certified correct and $n/4$ to be pure luck, refining ${\mathbb{E}}(X_{n})=\frac{3}{4}\cdot n$.
@@ 1.2. Notation
2. Shelf Shuffling
We begin by recalling the definition of the shelf shuffle with a single shelf.
@@ 2. Shelf Shuffling
Given an ordered deck of $n$ cards labelled one up to $n$. Cards are drawn from the bottom of the deck and placed in a pile or on a shelf. When a card is drawn, it is placed at the top or at the bottom of the pile with probability 1/2.
@@ 3.2. Distributional analysis
$S(z,v)=\frac{zv+\frac{\frac{z^{2}v}{2}}{1-\frac{zv}{2}}}{1-\frac{zv}{2}-\frac{\frac{z}{2}}{1-\frac{zv}{2}}+\frac{z}{2}}=\frac{2zv\big(2+(1-v)z\big)}{4-4vz+(v^{2}-v)z}.$
@@ 3.3. Expectation, variance and limit law
$\rho_{1}(v)=\frac{2(v-\sqrt{v})}{v(v-1)},\quad\rho_{2}(v)=\frac{2(v+\sqrt{v})}{v(v-1)},$
@@ 4.1. Asymmetric shelf shuffle
For $\frac{1}{2}\leq p<1$ we can carry out analysis as before, as the optimal strategy stays the same, but for $p<\frac{1}{2}$ the structure is asymptotically similar, but the optimal strategy changes, as we do not always guess the number one, as the guess depends on the value of $p$. Apparently, we have to study the equation $p=(1-p)^{n-1}$, allowing us to distinguish which label appear with the highest probability. Let $\nu=\lfloor\ln(p)/\ln(1-p)\rfloor+1$. Given a value of $n$ with $2\leq n\leq\nu$, we guess always the largest label $n$, but for all $n>\nu$ we guess the smallest label one. We note again that once card n1n-1 or $n$ is first shown, the player should guess the remaining cards in descending order, and is guaranteed to get these correct. After the first guess, we notice the label $j$ of the shown card. We now immediately know the positions of cards labelled one up to j1j-1. Concerning randomness we replace $n$ by njn-j and continue similar to starting with a set of njn-j cards by comparing njn-j to $\nu$ in order to make our choice for the next guess. Then, we continue the guessing, adding the number $j$ to our guess, as we actually have the set of labels $j+1,\dots,n$ left. We continue until n1n-1 or $n$ is shown and then correctly guess all the remaining cards.
Example 2 (Non-uniqueness of optimal strategy).
$p^{\ast}=\frac{1}{6}\left(2^{2/3}\sqrt[3]{\sqrt{93}-9}-2^{2/3}\sqrt[3]{9+\sqrt{93}}+6\right)\approx 0.31767<1/2$
@@
Consider the case of $n=4$ cards. For any $p<p^{\ast}$ we have to guess the label four as the first guess.
We summarize the strategies in the following result.
@@
For $p\in[\frac{1}{2},1]$ we proceed as in the case $p=\frac{1}{2}$: we guess the card labelled one as the first guess and then always the card immediately following the previously shown card in all other positions. Once one of card n1n-1 or $n$ is first shown, the player should guess the remaining cards in descending order, and is guaranteed to get these correct. For $p\in[0,\frac{1}{2})$ we guess the label $\operatorname{maxP}_{n}\in\{1,n\}$, corresponding to the maximum of of $\{p,(1-p)^{n-1}\}$. Let $\nu=\lfloor\ln(p)/\ln(1-p)\rfloor+1$. Then, we have
$\operatorname{maxP}_{n}=\begin{cases}n,\quad 2\leq n\leq\nu,\\ 1,\quad n>\nu.\end{cases}$
After the card, say labelled $j$ is shown to the guesser, we proceed in a similar way, but with a reduced set of indices $\{1,\dots,n-j\}$ and guess the label corresponding to the maximum of $\{p,(1-p)^{n-j}\}$, shifted of course by $j$.
@@
with initial values $X_{1}=1$ and $X_{0}=0$ and the random variable $\operatorname{FC}$ as stated in (1). Here, $\mathbb{I}_{n}(\operatorname{FC}_{n}=\operatorname{maxP}_{n})$ denotes the indicator of the event that the guess is correct and $X^{\ast}$ denotes an independent copy of $X$ and the random variables appearing in the decomposition are mutually independent.
As mentioned before, for $p\geq\frac{1}{2}$ we always have $\operatorname{maxP}_{n}=1$, and the distributional equation can be written as
with $I_{1}\operatorname{\overset{\mathcal{L}}{=}}\operatorname{Be}(p)$ denoting a Bernoulli distribution and initial values $X_{1}=1$ and $X_{0}=0$ and
The generating functions approach works directly exactly the same and we obtain for
$S(z,v)=S(z,v;p)=\sum_{n\geq 1}{\mathbb{E}}(v^{X_{n}})z^{n}$
the following result.
$S(z,v)=\frac{-z\big(1+(v-1)(p-1)z\big)v}{zv-1+pv(v-1)(p-1)z^{2}}$
We obtain by extraction of coefficients from $\partial_{v}S(z,1)$ and $\partial_{v}^{2}S(z,1)$ for the expected value and the variance for $n\geq 3$ and $\frac{1}{2}\leq p<1$ the results
${\mathbb{E}}(X_{n})$ $=(1-p+p^{2})n+3p-1-2p^{2},$
${\mathbb{V}}(X_{n})$ $=(1-p)p(3p^{2}-3p+1)n+p(1-p)(10p-8p^{2}-3).$
The denominator is readily factorized and we obtain the dominant singularity $\rho(v)$, depending on $\frac{1}{2}\leq p<1$
$\rho(v)=\frac{-v+\sqrt{(1+4p(p-1))(v^{2}-v)+v}}{2vp(pv-p-v+1)}.$
We can readily obtain a central limit theorem using the quasi-power theorem, such that
$\frac{X_{n}-\mu_{n}}{\sigma_{n}}\to\mathcal{N}(0,1),$
with $\mu_{n}={\mathbb{E}}(X_{n})$, $\sigma_{n}^{2}={\mathbb{V}}(X_{n})$; details are omitted. Of interest is also the transition at $p=1$. Here, we deterministically have $X_{n}=n$. Thus, we set $p=1-\frac{\lambda}{n^{\alpha}}$, with real $\alpha,\lambda>0$, which approaches one as $n$ tends to infinity. The transition can be easily observed by looking at the probability ${\mathbb{P}}\{X_{n}=n\}$, as it satisfies for $p=1-\frac{\lambda}{n^{\alpha}}$, $\alpha,\lambda>0$ and $n\to\infty$ the expansions
${\mathbb{P}}\{X_{n}=n\}\to\begin{cases}1,\quad 0<\alpha<1,\\ e^{-\lambda},\quad\alpha=1,\\ 0,\quad\alpha\geq 1,\end{cases}$
which follows directly from the explicit formula ${\mathbb{P}}\{X_{n}=n\}=p^{n}$ and the $\exp-\log$ expansion. More refined, we additionally anticipate here a phase transition for the shifted random variable $Z_{n}=n-X_{n}$ depending on the value of $\alpha$ from degenerated, to a discrete distribution, and then to a normal distribution. We expect a classical $\operatorname{Po}(\lambda)$ discrete limit in the case $\alpha=1$, motivated by the corresponding expansions of the expected value and the variance in this range
${\mathbb{E}}(X_{n})\sim n-c+\mathcal{O}\big(\frac{1}{n}\big),\quad{\mathbb{V}}(X_{n})\sim c+\mathcal{O}\big(\frac{1}{n}\big)$
This can be made precise by a closer inspection of the probability generating function using the dominant singularity; details are again omitted.
In the case of $0<p<\frac{1}{2}$ the analysis is more involved, depending on $\nu=\lfloor\ln(p)/\ln(1-p)\rfloor+1$, where the optimal strategy changes. First, we note that the distributional equation can be written as
$X_{n}\operatorname{\overset{\mathcal{L}}{=}}I_{1}X^{\ast}_{n-1}+(1-I_{1})\big(X^{\ast}_{n-J_{n}}+J_{n}-1+\mathbb{I}_{n}(J_{n}=n)\big),\quad n\leq\nu,$ (10)
and for $n\geq\nu$ by
$X_{n}\operatorname{\overset{\mathcal{L}}{=}}I_{1}\big(X^{\ast}_{n-1}+1\big)+(1-I_{1})(X^{\ast}_{n-J_{n}}+J_{n}-1).$
For the interested reader we sketch how to determine the distribution of $X_{n}$. In order to solve this, we introduce an auxiliary random variable $Y_{n}$, which satisfies (10) for all $n\geq 2$. Consequently,
[…]
Let $T=T(z,v)=\sum_{n\geq 1}{\mathbb{E}}(v^{Y_{n}})z^{n}$. We obtain the equation
$T-zv=zpT+\frac{zpT}{1-zv(1-p)}-zpT+\frac{z^{2}v^{2}(1-p)}{1vz(1-p)},$
which leads to a very simple explicit expression for $T$:
$T(z,v)=\frac{zv}{1-z\big(p+(1-p)v\big)},$
such that $Y_{n}$, and consequently also $X_{n}$ for $n\leq\nu$, is given by a sum of n1n-1 iid Bernoulli random variables $B_{k}\operatorname{\overset{\mathcal{L}}{=}}\operatorname{Be}(1-p)$ shifted by one:
$X_{n}\operatorname{\overset{\mathcal{L}}{=}}1+\sum_{k=1}^{n-1}B_{k}\operatorname{\overset{\mathcal{L}}{=}}1+\operatorname{Bin}(n-1,1-p),\quad 1\leq n\leq\nu.$
Then, we use again generating functions for treating (8). As before, let $S=S(z,v)=\sum_{n\geq 1}{\mathbb{E}}(v^{X_{n}})z^{n}$, which splits into two parts
$S(z,v)=S_{\leq\nu}(z,v)+S_{>\nu}(z,v)=\sum_{n=1}^{\nu}{\mathbb{E}}(v^{X_{n}})z^{n}+\sum_{n>\nu}{\mathbb{E}}(v^{X_{n}})z^{n},$ (11)
with $S_{\leq\nu}(z,v)=\sum_{n=1}^{\nu}{\mathbb{E}}(v^{X_{n}})z^{n}$ already known from $T(z,v)$:
$S_{\leq\nu}(z,v)=\sum_{n=1}^{\nu}{\mathbb{E}}(v^{Y_{n}})z^{n}=zv\frac{1-z^{\nu}\big(p+(1-p)v\big)^{\nu}}{1-z\big(p+(1-p)v\big)}.$ (12)
From (8) we obtain the equation
$S(z,v)-zv$ $=pz\big(S_{\leq\nu}(z,v)-(1-v)s_{\nu}(v)+vS_{>\nu}(z,v)\big)$
$\quad+\frac{zpS(z,v)}{1-zv(1-p)}-zpS(z,v)+\sum_{n\geq 2}(1-p)^{n-1}v^{n-1+\mathbb{I}_{n}(n\leq\nu)}z^{n}.$
This allows to characterize $S_{>\nu}(z,v)$, and thus the distribution of $X_{n}$ in terms of the $S_{\leq\nu}(z,v)$ (12) and $s_{\nu}(v)=v(p+v(1-p))^{\nu-1}$:
$S_{>\nu}(z,v)$ $=\frac{zv+(pz-1)S_{\leq\nu}(z,v)-p(1-v)zs_{\nu}(v)+\frac{zpS_{\leq\nu}(z,v)}{1-zv(1-p)}-zpS_{\leq\nu}(z,v)}{1-pzv-\frac{zp}{1-zv(1-p)}+zp}$
$\quad+\frac{1}{1-pzv-\frac{zp}{1-zv(1-p)}+zp}\sum_{n\geq 2}(1-p)^{n-1}v^{n-1+\mathbb{I}_{n}(n\leq\nu)}z^{n}.$
@@ 4.2. Refined analysis of the ordinary shelf shuffle
Here, one may apply the higher dimensional quasi-power theorem analogs of Heuberger and Kropf [14], leading to a central limit theorem, with variance-covariance governed by (11). We omit the details.
4.3. Refined analysis of the asymmetric shelf shuffle
Finally, we note that both ideas, asymmetric shelf shuffle and refined number of correct guesses, can be combined. For $\frac{1}{2}\leq p<1$ this leads to the distributional equation
for $n\geq 2$ with $L_{1}=L_{0}=0$, $C_{1}=1$ and $C_{0}=0$, and $I_{1}=\operatorname{Be}(p)$. We only state the corresponding generating function $S(z,v,w)=S(z,v,w,p)$:
$S(z,v,w)=\frac{zv\big(1+(v-1)(-1+p)z\big)}{1-pv(w-1)(-1+p)z^{2}-(1-p)v+wpz},$
from which results for the expected values, variances and the covariance can readily be obtained, as well as distributional results. It is also possible to extend this analysis to the case $0<p<\frac{1}{2}$.
@@ 5. Outlook and summary
A corresponding analysis of the guessing game without feedback seems to be more involved, insofar as even the optimal strategy is unclear. A spectral analysis of the position matrix for the asymmetric model is doable. Of interest is also the transition for $p\to 0$ in the asymmetric model and to look again for transitions from normal to say Poisson or degenerated.